OpenStudy (anonymous):

how do i differentiate abs(x^2-1)? please help me

OpenStudy (dumbcow):

you have to split it up because it cannot be differentiated at points x= +-1

OpenStudy (mrdoe):

rewrite as \[\sqrt{(x ^{2}-1}^{2}\]

OpenStudy (anonymous):

work in cases

OpenStudy (mrdoe):

except with correct placement on that parentheses

OpenStudy (anonymous):

if \(x<-1\) or \(x>1\) it is \(x^2-1\) but if \(-1<x<1\) it is \(1-x^2\)

OpenStudy (anonymous):

both of those are very easy to differentiate your derivative is a piecewise function

OpenStudy (anonymous):

which is not surprising since so is \(f(x)=|x^2-1|\)

OpenStudy (anonymous):

so I should differentiate x^2-1 and 1-2x^2?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

well no not \(1-2x^2\)just \(1-x^2\)

OpenStudy (anonymous):

so 2x and -2x are the answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

depending on the values of \(x\) it is piecewise

OpenStudy (anonymous):

you are awesome bro can i adopt you?

OpenStudy (anonymous):

\[f'(x) = \left\{\begin{array}{rcc} 2x& \text{if} & x< -1 \text{ or} x>1\\ -2x& \text{if} &-1< x < 1 \end{array} \right. \]

OpenStudy (anonymous):

which is exactly because \[f(x) = |x^2-1| = \left\{\begin{array}{rcc} x^2-1 & \text{if} & x<-1 \text { or } x>1 \\ 1-x^2& \text{if} & -1<x < 1 \end{array} \right. \]

OpenStudy (anonymous):

thanks you can buy me a whiskey at the bar

OpenStudy (anonymous):

thank you.I'll send you that whiskey asap :D